# Performance Analysis of NR Sidelink and Wi-Fi Coexistence Networks in Unlicensed Spectrum

### v1

@@ Abstract
With the rapid development of various internet of things (IoT) applications, including industrial IoT (IIoT) and visual IoT (VIoT), the demand for direct device-to-device communication to support high data rates continues to grow. To address this demand, 5G-Advanced has introduced sidelink communication over the unlicensed spectrum (SL-U) as a method to increase data rates. However, the primary challenge of SL-U in the unlicensed spectrum is ensuring fair coexistence with other incumbent systems, such as Wi-Fi. In this paper, we address the challenge by designing channel access mechanisms and power control strategies to mitigate interference and ensure fair coexistence. First, we propose a novel collaborative channel access (CCHA) mechanism that integrates channel access with resource allocation through collaborative interactions between base stations (BS) and SL-U users. This mechanism ensures fair coexistence with incumbent systems while improving resource utilization. Second, we mathematically model the joint channel access and power control problems, analyzing the trade-off between fairness and transmission rate to minimize interference and optimize performance in the coexistence system. Finally, we develop a collaborative subgoal-based hierarchical deep reinforcement learning (C-GHDRL) framework. This framework enables SL-U users to make globally optimal decisions by leveraging collaborative operations between the BS and SL-U users, effectively overcoming the limitations of traditional optimization methods in solving joint optimization problems with nonlinear constraints. Simulation results demonstrate that the proposed scheme significantly enhances the coexistence system’s performance while ensuring fair coexistence between SL-U and Wi-Fi users.
@@ I Introduction
With the advent of industry 5.0, driven by emerging applications such as predictive maintenance, hyper-customization in manufacturing, and collaborative robotics [1], the data rates of future Industrial Internet of Things (IIoT) systems are predicted to increase by 10 to 50 times compared to current networks [2]. Additionally, the Visual Internet of Things (VIoT) plays a crucial role in urban counter-terrorism, intelligent border security, and missing person searches by leveraging visual sensors to process and analyze real-time visual data streams intelligently [3]. Either the predictive maintenance within IIoT or the transmission of video streams in VIoT, sidelink (SL) communication with high data rates is required to deliver these services effectively. To address the demand, the evolution of new radio (NR) sidelink [4], one of the initial projects approved by 5G-Advanced, offers a solution by utilizing sidelink over unlicensed spectrum (SL-U). SL-U is a novel wireless communication protocol that enables devices to communicate directly over unlicensed spectrum without relying on cellular networks or access points. It is highly suitable for short-distance wireless communication applications like the Internet of Things (IoT) [5].
@@
While the introduction of SL-U can significantly enhance the data rate of wireless links, the primary challenge for SL-U is to ensure fair coexistence with other incumbent systems, such as Wi-Fi, in the unlicensed spectrum. The most critical of fair coexistence is the need for effective interference mitigation to ensure fair and efficient spectrum sharing [6, 7, 8]. To address the challenge, SL-U employs the listen-before-talk (LBT) channel access mechanism, which enables devices to perform a clear channel assessment (CCA) to sense channel availability prior to transmission [9]. However, the 3rd Generation Partnership Project (3GPP) standards specified that the minimum scheduling resource of physical sidelink shared channel (PSSCH) in the time domain is a slot [10]. Consequently, after completing the LBT process, SL-U users in the unlicensed spectrum are required to wait for the spectrum slot boundary (SSB) before initiating data transmission. This requirement may lead to channel access collisions and inefficient resource utilization. To address this issue, [11] proposed a method to reduce collisions by randomizing the starting position of the back-off countdown process within a slot. However, the approach is limited to SL-U-only networks and does not account for coexistence scenarios. [12] proposed a new CR-LBT (LBT With Collision Resolution) method of transmitting reserved signals, which greatly reduces collisions and enhances the fairness of channel resource sharing. However, the presence of reserved signals wastes channel resources.
@@
In addition, selecting the appropriate transmission power for SL-U users after channel access is also essential for interference mitigation [13, 14, 15]. Considering inter-cell and inter-operator interference in both licensed and unlicensed spectrum, [16] proposed a joint optimization framework that integrates user association, power allocation, and dynamic spectrum sharing to maximize network throughput. [17] presented a joint power control and time allocation scheme for licensed and unlicensed spectrum in unmanned aerial vehicle (UAV)-assisted Internet of Vehicles (IoV) systems, optimizing power and time allocation to maximize overall system capacity.
@@
Channel access and power control are two fundamental mechanisms for SL-U in the coexistence network, and they exhibit significant interdependence [11, 13, 18]. This interdependence becomes especially significant in unlicensed spectrum coexistence scenarios. For example, considering the coexistence of terrestrial cellular networks, Wi-Fi, and UAV-based base stations (UAV-BSs), [19] jointly optimized sub-channel allocation and power control of UAV users across the licensed and unlicensed spectrum to maximize the total uplink rate. [20] jointly optimized the time and power allocation during maximum channel occupation time (MCOT) to maximize the total throughput of both downlink and uplink in New Radio on unlicensed channel spectrum (NR-U) while ensuring fair coexistence with Wi-Fi. To this end, we jointly optimize channel access and power control to mitigate interference and enhance spectrum utilization. Various centralized algorithms based on combinatorial optimization [21], random geometry [22], and game theory [23] have been proposed for the joint optimization problem. However, the joint optimization of channel access and power control is often formulated as a combinatorial optimization problem with nonlinear constraints, which leads to the ineffectiveness of traditional optimization methods in most scenarios [13, 24].
@@
The emergence of deep reinforcement learning (DRL) has introduced promising solutions to joint optimization problems with nonlinear constraints [25, 26]. In [27], a distributed DRL-based scheme was proposed, enabling D2D pairs to autonomously optimize channel selection and transmission power using only local information and outdated non-local information. [28] introduced a novel DRL-based framework for jointly optimizing power and block length allocation to minimize the worst-case decoding-error probability in the finite block length (FBL) regime for an ultra-reliable low-latency communication (URLLC) -based downlink V2X communication system. However, it is important to note that DRL-based approaches typically require all actions to operate on the same time scale when solving joint optimization problems, meaning that all actions must be decided simultaneously. This assumption may not be applicable in certain communication scenarios where actions occur at different timescales.
@@
Therefore, hierarchical deep reinforcement learning (HDRL) is proposed to solve the problem [29, 30, 31, 32]. By adopting a hierarchical learning framework, HDRL can achieve both global and local optimization, which is very suitable for the joint optimization of channel access and power control with different timescales. In [33], the authors optimized the CCA threshold and transmit power of the access point (AP) using the HDRL algorithm, improving network performance. However, [33] focuses solely on a scenario involving Wi-Fi users and does not address the coexistence challenges associated with the unlicensed spectrum. [34] proposed leveraging the HDRL algorithm to jointly optimize partial spectrum sensing and power allocation, enabling users to make intelligent decisions on spectrum selection and power allocation to maximize network throughput while minimizing mutual interference. Although each secondary user (SU) transmits information to the central server to train the HDRL neural network, the central server maintains a separate model for each user. As a result, each network is trained based solely on the local information available to the corresponding user, which may prevent the model from achieving a globally optimal solution.
@@
In summary, to address the issues of channel collisions and low resource utilization caused by time slot alignment for SL-U users, we propose a novel collaborative channel access (CCHA) mechanism. Additionally, we model the joint channel access and power control problem, analyzing strategies to enhance the overall performance of the coexistence system while balancing the trade-off between fairness and transmission rate. Finally, to overcome the challenges posed by the lack of an information-sharing mechanism between SL-U users in the unlicensed spectrum, which prevents models from making globally optimal decisions, and the inefficiency of traditional combinatorial optimization methods under nonlinear constraints, we developed a collaborative subgoal-based HDRL (C-GHDRL) algorithm framework. The framework achieves joint optimization of channel access and power control through collaborative interactions between the BS and SL-U users, thereby significantly improving the overall performance of the coexistence system. The main contributions of this paper are as follows:
@@
  • We propose a CCHA mechanism integrating channel access and resource allocation. In this mechanism, the BS allocates available resources to SL-U users within its coverage area based on channel sensing results. SL-U users assigned to these resources perform lightweight LBT before accessing the channel. This collaborative interaction between the BS and SL-U users not only ensures complete resource orthogonality but also increases the opportunities for channel access for SL-U users. Consequently, this mechanism enhances resource utilization and significantly improves coexistence system performance.
  • For the SL-U and Wi-Fi coexistence system, we establish a mathematical joint model for the channel access and power control problems to analyze how to balance fairness and transmission rate under relevant constraints.
@@
  • We propose a C-GHDRL algorithm framework designed to jointly optimize channel access and power control for SL-U users in coexistence systems. In the framework, BS collects global information through interaction with SL-U users to train the C-GHDRL model and periodically distributes the trained model to all SL-U users. Each SL-U user applies this model autonomously to determine optimal channel access and transmission power. Through collaborative operations between the BS and SL-U users, the algorithm enables SL-U users to make globally optimal decisions while effectively addressing the joint optimization of channel access and power control under nonlinear constraints, thereby significantly improving the overall performance of coexistence systems.
@@ II System Model and Problem Formulation
Consider a communication scenario in which Wi-Fi and SL-U coexist in the unlicensed spectrum, as depicted in Fig. 1. This scenario includes a Wi-Fi AP, an SL-U BS, multiple Wi-Fi users, and multiple SL-U user pairs. Each SL-U user pair comprises a transmitter (Tx) and a receiver (Rx). The BS and AP are centrally located within the area, while Wi-Fi and SL-U users are randomly distributed. The distance between the Tx and Rx of each SL-U user pair is represented by a random variable, denoted as $dis_{random}$. Let =$\mathcal{M}=\{1,2,\cdots,M\}$ denote the set of SL-U user pairs, and $\mathcal{N}=\{1,2,\cdots,N\}$ denote the set of Wi-Fi users. Define $T_{m}$ and $R_{m}$ as the transmitter and receiver of SL-U user pair $m\in\mathcal{M}$ , respectively.
@@
    LBT: The LBT scheme adopts a similar back-off mechanism to that of the CSMA/CA protocol. Before initiating data transmission, an SL-U user performs CCA to sense whether the channel is busy or idle. If the channel remains idle for a specified period, denoted as $T_{d}$, the user initiates a back-off process, introducing a random time before attempting transmission. However, unlike in CSMA/CA, where users can transmit immediately after the back-off process concludes, SL-U users must wait for the next SSB before transmitting data. [9] defined two types of channel access processes for LBT:
    • Type 1: Channel access procedure with random back-off of variable size contention window (CW). According to [9], the CW size is determined by the channel access priority $p$, as indicated in Table I.
    • Type 2: No random backoff procedure is carried out, and only a predetermined period of channel sensing or no channel sensing is performed before accessing the channel.
@@
For the Wi-Fi users path loss model, the free space loss model is used:
$PL_{w}=32.44+20\times\log_{10}{f_{c}}+20\times\log_{10}{(d_{w})},$ (1)
where $f_{c}$ and $d_{w}$ are the center frequency in GHz and the distance between the Wi-Fi user and the AP in meters, respectively.
The SL-U users path loss model is based on the 3GPP standard and applies to line-of-sight (LOS) office scenarios[36]. The path loss model is as follows:
@@
where $d_{s}$ is the distance between the transmitter and receiver of the SL-U user in meters. The corresponding shadow fading (SF) follows a log-normal distribution with a mean of 0 and a standard deviation of $\sigma_{SF}=3$.
Therefore, the signal-to-interference-to-noise ratio (SINR) of the SL-U pair $m$ and the Wi-Fi user $n$ in time slot $t$ are respectively defined as:
$\eta_{s,m}(t)=\frac{p_{s,m}^{T}(t)-PL_{s,m}-PL_{SF,m}}{p_{noise}+\sum\limits_{n=0}^{N}{k_{w,n}(t)p_{inter,w}^{R_{m},n}(t)}},$ (3)
@@
where $p_{s,m}^{T}(t)$ and $p_{w,n}^{T}(t)$ are the transmission power of SL-U user $T_{m}$ and the Wi-Fi user $n$, respectively. The $PL_{s,m}$ and $PL_{SF,m}$ represent respectively the path loss and the shadow fading between the transmitter and receiver of the SL-U pair $m$. The $PL_{w,n}$ is the path loss of the Wi-Fi user $n$. The $p_{noise}$ is additive Gaussian white noise (AWGN) power, $p_{inter,w}^{R_{m},n}(t)$ represents the interference power received by SL-U $R_{m}$ from Wi-Fi user $n$, $p_{inter,s}^{AP,m}(t)$ represents the interference power received by AP from SL-U $T_{m}$. If $k_{s,m}(t)=1$ indicates that SL-U pair $m$ accesses the channel during time slot $t$; otherwise, $k_{s,m}(t)=0$ indicates that SL-U pair $m$ does not access the channel. If $k_{w,n}(t)=1$ indicates that Wi-Fi user $n$ accesses the channel during time slot $t$; otherwise, $k_{w,n}(t)=0$ indicates that Wi-Fi user $n$ does not access the channel.
@@
Fairness measure is used in network engineering to determine whether users or applications are receiving a fair share of system resources. The formula for Jain’s fairness index is as follows:
@@
where $x_{i}$ is the throughput for the $i$th connection. $n$ denotes the total number of users. According to the characteristics of Jain’s fairness index, it is a continuous value from $1/n$ to $1$. The index is $1$ when all users receive the same allocation[37].
@@
$\ s.t.&C1:p_{s,min}^{T}\leq p_{s,m}^{T}(t)\leq p_{s,max}^{T},\forall m\in\mathcal{M},\\ &\\ &C2:p_{w,min}^{T}\leq p_{w,n}^{T}(t)\leq p_{w,max}^{T},\forall n\in\mathcal{N},\\ &\\ &C3:\sum\limits_{m=0}^{M}{k_{s,m}(t)}\leq 1,\forall m\in\mathcal{M},\\ &\\ &C4:\sum\limits_{n=0}^{N}{k_{w,n}(t)}\leq 1,\forall n\in\mathcal{N}.\\ \end{array}$
where constraints $C1$ and $C2$ restrict the transmission power of each SL-U user pair and Wi-Fi user, respectively. Constraints $C3$ indicates that only one SL-U user pair occupies the channel per slot at most. Constraint $C4$ shows that at most one Wi-Fi user occupies the channel per slot.

### v2

# Collaborative Channel Access and Transmission for NR Sidelink and Wi-Fi Coexistence over Unlicensed Spectrum

@@ Abstract
With the rapid development of various internet of things (IoT) applications, including industrial IoT (IIoT) and visual IoT (VIoT), the demand for direct device-to-device communication to support high data rates continues to grow. To address this demand, 5G-Advanced has introduced sidelink communication over the unlicensed spectrum (SL-U) to increase data rates. However, the primary challenge of SL-U in the unlicensed spectrum is ensuring fair coexistence with other incumbent systems, such as Wi-Fi. In this paper, we address the challenge by designing channel access mechanisms and power control strategies to mitigate interference and ensure fair coexistence. First, we propose a novel collaborative channel access (CCHA) mechanism that integrates channel access with resource allocation through collaborative interactions between base stations (BS) and SL-U users. This mechanism ensures fair coexistence with incumbent systems while improving resource utilization. Second, to further enhance the performance of the coexistence system, we develop a cooperative subgoal-based hierarchical deep reinforcement learning (C-GHDRL) algorithm framework. The framework enables SL-U users to make globally optimal decisions by leveraging cooperative operations between the BS and SL-U users, effectively overcoming the limitations of traditional optimization methods in solving joint optimization problems with nonlinear constraints. Finally, we mathematically model the joint channel access and power control problem and balance the trade-off between fairness and transmission rate in the coexistence system by defining a suitable reward function in the C-GHDRL algorithm. Simulation results demonstrate that the proposed scheme significantly enhances the performance of the coexistence system while ensuring fair coexistence between SL-U and Wi-Fi users.
@@ I Introduction
With the advent of industry 5.0, driven by emerging applications such as predictive maintenance, hyper-customization in manufacturing, and collaborative robotics [1], the data rates of future Industrial Internet of Things (IIoT) systems are predicted to increase by 10 to 50 times compared to current networks [2]. Additionally, the Visual Internet of Things (VIoT) plays a crucial role in urban counter-terrorism, intelligent border security, and missing person searches by leveraging visual sensors to process and analyze real-time visual data streams intelligently [3]. Either the predictive maintenance within IIoT or the transmission of video streams in VIoT, sidelink (SL) communication with high data rates is required to deliver these services effectively. To address the demand, the evolution of new radio (NR) sidelink [4], one of the initial projects approved by 5G-Advanced, offers a solution by utilizing sidelink over unlicensed spectrum (SL-U). SL-U combines sidelink and NR on unlicensed channel spectrum (NR-U) technologies [5][6], leveraging idle unlicensed spectrum to enhance the data rates. So, it is highly suitable for short-distance wireless communication applications like the Internet of Things (IoT) [7].
@@
While the introduction of SL-U can significantly enhance the data rate of wireless links, the primary challenge for SL-U is to ensure fair coexistence with other incumbent systems, such as Wi-Fi, in the unlicensed spectrum. The most critical of fair coexistence is the need for effective interference mitigation to ensure fair and efficient spectrum sharing [8][9]. To address the challenge, SL-U employs the listen-before-talk (LBT) channel access mechanism, which enables devices to perform a clear channel assessment (CCA) to sense channel availability prior to transmission [10]. Although the LBT mechanism effectively mitigates interference by selecting idle channels, the existing LBT channel access mechanism may lead to channel collisions and low resource utilization for SL-U users due to slot alignment [6][11]. To address this issue, [6] proposed a method to reduce collisions by randomizing the starting position of the back-off countdown process within a slot. However, the approach does not account for coexistence scenarios. [11] proposed a new CR-LBT (LBT With Collision Resolution) method of transmitting reserved signals, which greatly reduces collisions and enhances the fairness of channel resource sharing. However, the presence of reserved signals wastes channel resources.
@@
Although the LBT mechanism effectively mitigates interference, efficient resource reuse in specific coexistence scenarios can further enhance spectrum utilization. During the process, power control is essential for effectively mitigating interference [12][13][14][15]. Considering inter-cell and inter-operator interference in both licensed and unlicensed spectrum, [14] proposed a joint optimization framework that integrates user association, power allocation, and dynamic spectrum sharing to maximize network throughput. [15] presented a joint power control and time allocation scheme for licensed and unlicensed spectrum in unmanned aerial vehicle (UAV)-assisted Internet of Vehicles (IoV) systems, optimizing power and time allocation to maximize overall system capacity.
@@
Therefore, channel access and power control are two fundamental mechanisms for SL-U in the coexistence network, and they exhibit significant interdependence [6][12][16]. This interdependence becomes especially significant in unlicensed spectrum coexistence scenarios. For example, considering the coexistence of terrestrial cellular networks, Wi-Fi, and UAV-based base stations (UAV-BSs), [17] jointly optimized sub-channel allocation and power control of UAV users across the licensed and unlicensed spectrum to maximize the total uplink rate. [18] jointly optimized the time and power allocation during maximum channel occupation time (MCOT) to maximize the total throughput of both downlink and uplink in NR-U while ensuring fair coexistence with Wi-Fi. Various centralized algorithms based on combinatorial optimization [19], random geometry [20], and game theory [21] have been proposed for the joint optimization problem. However, the joint optimization of channel access and power control is often formulated as a combinatorial optimization problem with nonlinear constraints, which leads to the ineffectiveness of traditional optimization methods in most scenarios [12][22].
@@
The emergence of deep reinforcement learning (DRL) has introduced promising solutions to joint optimization problems with nonlinear constraints [23][24][25][26]. In [25], a distributed DRL-based scheme was proposed, enabling D2D pairs to autonomously optimize channel selection and transmission power using only local information and outdated non-local information. [26] introduced a novel DRL-based framework for jointly optimizing power and block length allocation to minimize the worst-case decoding-error probability in the finite block length (FBL) regime for an ultra-reliable low-latency communication (URLLC) -based downlink V2X communication system. However, it is important to note that DRL-based approaches typically require all actions to operate on the same time scale when solving joint optimization problems, meaning that all actions must be decided simultaneously. This assumption may not be applicable in certain communication scenarios where actions occur at different timescales.
@@
Therefore, hierarchical deep reinforcement learning (HDRL) is proposed to solve the problem of joint optimization that typically requires all operations to be at the same time scale [27][28][29][30][31]. By adopting a hierarchical learning framework, HDRL can achieve both global and local optimization, which is very suitable for the joint optimization of channel access and power control with different timescales. In [30], the authors optimized the CCA threshold and transmit power of the access point (AP) using the HDRL algorithm, improving network performance. However, [30] focuses solely on a scenario involving Wi-Fi users and does not address the coexistence challenges associated with the unlicensed spectrum. [31] proposed leveraging the HDRL algorithm to jointly optimize partial spectrum sensing and power allocation, enabling users to make intelligent decisions on spectrum selection and power allocation to maximize network throughput while minimizing mutual interference. Although each secondary user (SU) transmits information to the central server to train the HDRL neural network, the central server maintains a separate model for each user. As a result, each network is trained based solely on the local information available to the corresponding user, which may prevent the model from achieving a globally optimal solution.
@@
In summary, we propose a novel collaborative channel access (CCHA) mechanism to address channel collisions and low resource utilization for SL-U users. Although the mechanism can improve resource utilization, effective resource reuse can further enhance resource utilization in some coexistence scenarios. Therefore, we jointly optimize channel access and power control to maximize spectrum efficiency. To overcome the inefficiencies of traditional combinatorial optimization methods under nonlinear constraints and joint optimization across different time scales, we developed a collaborative subgoal-based HDRL (C-GHDRL) algorithm framework. The framework achieves joint optimization of channel access and power control through collaborative interactions between the BS and SL-U users, thereby significantly improving the overall performance of the coexistence system. Finally, we model the joint channel access and power control problem and achieve a balance between fairness and transmission rate by defining an appropriate reward function within the algorithm. The main contributions of this paper are as follows:
@@
  • We propose a CCHA mechanism integrating channel access and resource allocation to ensure fair coexistence. In this mechanism, the BS allocates available resources to SL-U users within its coverage area based on channel sensing results. SL-U users assigned to these resources perform lightweight LBT before accessing the channel. This collaborative interaction between the BS and SL-U users not only ensures complete resource orthogonality but also increases the opportunities for channel access for SL-U users. Consequently, this mechanism enhances resource utilization and significantly improves coexistence system performance.
@@
  • To further enhance the coexistence system performance, we propose a C-GHDRL algorithm framework designed to jointly optimize channel access and power control for SL-U users in coexistence systems. In the framework, BS collects global information through interaction with SL-U users to train the C-GHDRL model and periodically distributes the trained model to all SL-U users. Each SL-U user applies this model autonomously to determine optimal channel access and transmission power. Through collaborative operations between the BS and SL-U users, the algorithm enables SL-U users to make globally optimal decisions while effectively addressing the joint optimization of channel access and power control under nonlinear constraints, thereby significantly improving the overall performance of coexistence systems.
  • We formulate a mathematical joint model for the channel access and power control problems as the joint optimization goal. By defining an appropriate reward function within the C-GHDRL algorithm, we balance fairness and transmission rate in the coexistence system.
@@ II System Model and Problem Formulation
Consider a communication scenario in which Wi-Fi and SL-U coexist in the unlicensed spectrum, as depicted in Fig. 1. This scenario includes a Wi-Fi AP, an SL-U BS, multiple Wi-Fi users, and multiple SL-U user pairs. Each SL-U user pair comprises a transmitter (Tx) and a receiver (Rx). The BS and AP are centrally located within the area, while Wi-Fi and SL-U users are randomly distributed. Let =$\mathcal{M}=\{1,2,\cdots,M\}$ denote the set of SL-U user pairs, and $\mathcal{N}=\{1,2,\cdots,N\}$ denote the set of Wi-Fi users. Define $T_{m}$ and $R_{m}$ as the transmitter and receiver of SL-U user pair $m\in\mathcal{M}$ , respectively.
@@
    LBT: The LBT scheme adopts a similar back-off mechanism to that of the CSMA/CA protocol. Before initiating data transmission, an SL-U user performs CCA to sense whether the channel is busy or idle. If the channel remains idle for a specified period, denoted as $T_{d}$, the user initiates a back-off process, introducing a random time before attempting transmission. [10] defined two types of channel access processes for LBT:
    • Type 1: Channel access procedure with random back-off of variable size contention window (CW). According to [10], the CW size is determined by the channel access priority $p$, as indicated in Table I. For example, when $p=1$, the time a user requires to sense the channel is given by $T_{sensing}=T_{d}+m_{p}\times T_{sl}+CW_{p}\times T_{sl}=34\mu s+CW_{p}\times 9\mu s$, where $T_{sl}$ is sensing slot.
[…]
@@
The Wi-Fi users and SL-U users path loss model is based on the 3GPP standard and applies to line-of-sight (LOS) scenarios[33]. The path loss model is as follows:
$PL_{w}=32.4+20\times\log_{10}{f_{c}}+17.3\times\log_{10}{(d_{w})},$ (1)
@@
where $f_{c}$ is the center frequency in GHz.The $d_{w}$ is the distance between the Wi-Fi user and the AP in meters, and the $d_{s}$ is the distance between the transmitter and receiver of the SL-U user in meters.
Therefore, the signal-to-interference-to-noise ratio (SINR) of the SL-U pair $m$ and the Wi-Fi user $n$ in slot $t$ are respectively defined as:
$\eta_{s,m}(t)=\frac{p_{s,m}^{T}(t)-PL_{s,m}}{p_{noise}+\sum\limits_{n=0}^{N}{k_{w,n}(t)p_{inter,w}^{R_{m},n}(t)}},$ (3)
@@
where $p_{s,m}^{T}(t)$ and $p_{w,n}^{T}(t)$ are the transmission power of SL-U user $T_{m}$ and the Wi-Fi user $n$, respectively. The $PL_{s,m}$ represents the path loss between the transmitter and receiver of the SL-U pair $m$. The $PL_{w,n}$ is the path loss of the Wi-Fi user $n$. The $p_{noise}$ is additive Gaussian white noise (AWGN) power, $p_{inter,w}^{R_{m},n}(t)$ represents the interference power received by SL-U $R_{m}$ from Wi-Fi user $n$, $p_{inter,s}^{AP,m}(t)$ represents the interference power received by AP from SL-U $T_{m}$. If $k_{s,m}(t)=1$ indicates that SL-U pair $m$ accesses the channel during slot $t$; otherwise, $k_{s,m}(t)=0$ indicates that SL-U pair $m$ does not access the channel. If $k_{w,n}(t)=1$ indicates that Wi-Fi user $n$ accesses the channel during slot $t$; otherwise, $k_{w,n}(t)=0$ indicates that Wi-Fi user $n$ does not access the channel.
@@
Fairness measure is used in network engineering to determine whether users or applications are receiving a fair share of system resources. The formula for Jain’s fairness index is as follows [34]:
@@
where $x_{i}$ is the throughput for the $i$th connection. $n$ denotes the total number of users. According to the characteristics of Jain’s fairness index, it is a continuous value from $1/n$ to $1$. The index is $1$ when all users receive the same allocation.
@@
$\ s.t.&C1:p_{s,min}^{T}\leq p_{s,m}^{T}(t)\leq p_{s,max}^{T},\forall m\in\mathcal{M},\\ &\\ &C2:p_{w,min}^{T}\leq p_{w,n}^{T}(t)\leq p_{w,max}^{T},\forall n\in\mathcal{N},\\ &\\ &C3:\sum\limits_{m=0}^{M}{k_{s,m}(t)}\leq 1,\forall m\in\mathcal{M},\\ &\\ &C4:\sum\limits_{n=0}^{N}{k_{w,n}(t)}\leq 1,\forall n\in\mathcal{N}.\\ &\\ &C5:\sum\limits_{m=0}^{M}{k_{s,m}(t)}+\sum\limits_{n=0}^{N}{k_{w,n}(t)}\leq 2,\forall m\in\mathcal{M},n\in\mathcal{N}.\\ \end{array}$
where constraints $C1$ and $C2$ restrict the transmission power of each SL-U user pair and Wi-Fi user, respectively. Constraints $C3$ indicates that only one SL-U user pair occupies the channel per slot at most. Constraint $C4$ shows that at most one Wi-Fi user occupies the channel per slot. Constraint $C5$ indicates that at most one Wi-Fi user and one SL-U user can use the channel simultaneously per slot.